OpenStudy (anonymous):

1. If x+3/3=y+2/2, then x/3 = ____. A. y+1 B. y/3 C. y/2 D. y-1 2.A model is made of a car. The car is 10 feet long and the model is 7 inches long. What is the ratio of the length of the car to the length of the model? A. 10 : 7 B. 7 : 120 C. 120 : 7 D. 7 : 10 3. 4/9=m/54 A. 24 B. 1/24 C. 6 D. 2/3

OpenStudy (anonymous):

x+3/3=y+2/2 the left side needs to look like x/3 First we want to isolate x on the left side. What should we do?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Do you know how to multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Multiply both sides by 3

OpenStudy (anonymous):

9

OpenStudy (anonymous):

lol you don't know how to multiply?

OpenStudy (anonymous):

Well let's do an easier one first # 3 4/9=m/54 \[\frac{4}{9}=\frac{m}{54}\]

OpenStudy (anonymous):

Multiply both sides by 54 in order to isolate m on the right side. Can you tell me what you get for m?

OpenStudy (anonymous):

so 54 *4 and 9

OpenStudy (anonymous):

yes \[\frac{54*4}{9}=m\] can you simplify that?

OpenStudy (anonymous):

486/9=m

OpenStudy (anonymous):

54=m

OpenStudy (anonymous):

How did you get 486?

OpenStudy (anonymous):

54*4

OpenStudy (anonymous):

No that doesn't sound right. Do it again.

OpenStudy (anonymous):

216